OpenStudy (waheguru):

It takes a boat 3 hours to travel down a river from point A to point B, and 5 hours to travel up the river from B to A. How long would it take the same boat to go from A to B in still water? How Do I solve this

OpenStudy (waheguru):

So...

OpenStudy (waheguru):

no one?

OpenStudy (anonymous):

maybe like this: let boat velocity be B and river velocity W 3(B+W)=5(B-W) (B-W)/(B+W)=3/5 so: B-W = 3 B+W = 5 from here B=4, W=1 distance between AB: 3(B+W)=3*5=15 so it would take: 4*t =15 t=15/4 hours

OpenStudy (waheguru):

CAn you expliain more

OpenStudy (anonymous):

what you don't understand?

OpenStudy (anonymous):

anything

OpenStudy (anonymous):

how do get "3(B+W)=5(B-W)"?